# PoseBench3D: A Cross-Dataset Analysis Framework for 3D Human Pose Estimation

### v1

@@ Abstract
Reliable three-dimensional human pose estimation is becoming increasingly important for real-world applications, yet much of prior work has focused solely on the performance within a single dataset. In practice, however, systems must adapt to diverse viewpoints, environments, and camera setups—conditions that differ significantly from those encountered during training, which is often the case in real-world scenarios. To address these challenges, we present a standardized testing environment in which each method is evaluated on a variety of datasets, ensuring consistent and fair cross-dataset comparisons—allowing for the analysis of methods on previously unseen data. Therefore, we propose PoseBench3D, a unified framework designed to systematically re-evaluate prior and future models across four of the most widely used datasets for human pose estimation—with the framework able to support novel and future datasets as the field progresses. Through a unified interface, our framework provides datasets in a pre-configured yet easily modifiable format, ensuring compatibility with diverse model architectures. We re-evaluated the work of 18 methods, either trained or gathered from existing literature, and reported results using both Mean Per Joint Position Error (MPJPE) and Procrustes Aligned Mean Per Joint Position Error (PA-MPJPE) metrics, yielding more than 100 novel cross-dataset evaluation results. Additionally, we analyze performance differences resulting from various pre-processing techniques and dataset preparation parameters—offering further insight into model generalization capabilities.
@@ 1 Introduction
Three-dimensional Human Pose Estimation (HPE) has gathered substantial interest for it’s critical role in applications such as healthcare [36, 29], action recognition [17, 16], military operations [46, 12], human-computer interaction [30, 38], and virtual/augmented reality [23]—among many others. Despite remarkable progress in recent years [14, 31, 47, 28, 9, 33], most work focuses on performance within a single, controlled dataset. The majority of existing work adopts either a direct-from-image approach or a two-stage approach, with the latter being widely used for its flexibility. In the two-stage approach, the first stage detects 2D keypoints, and the second stage, known as the “lifting network” [22, 42, 15, 4, 49, 11, 40] maps the 2D keypoints to 3D pose estimations. However, these two-stage models show poor generalization across various different datasets [35, 20].
@@
To assess the generalization of these models, we consider cross-dataset performance: the task of training on one dataset and testing on another. Machine learning models are generally susceptible to overfitting to the bias and variance of the data distribution it was trained on; this is particularly pronounced in 3D human pose estimation datasets due to variations in data collection procedures, where the underlying distributions of datasets often differ significantly. For example, the H36M dataset [10] was captured using a four-camera setup, while GPA [34] used five, and both 3DPW [21] and SURREAL [32] relied on a single-camera configuration. In fig. 1, we show the frequency histogram of camera positions relative to the subject in spherical coordinates. While elevation distributions are largely consistent across datasets (cf. fig. 1(a)), the azimuthal angles vary significantly (cf. fig. 1(b)).
Beyond camera setup, joint placement conventions also differ between datasets. For instance, there is currently no standardized convention for placing the hip or spine joints (cf. fig. 2)—note the difference in width of hips in fig. 2(a) and fig. 2(c). Additionally, since datasets are collected with varying action sets, the captured range of motion also varies. In fig. 3, the blue dots represent the recorded range of motion, whereas the yellow region shows the valid range. Since these two datasets do not explore the entire valid range of motion, normalization performed with one will not translate well onto the other.
Research Gap. These vast differences mean that, although any method may achieve near-optimal performance within its own dataset, it is likely to perform poorly across differing datasets. Measuring cross-dataset performance is paramount, but is extremely labor-intensive to perform manually. Our goal in this paper is to automate the task of cross-dataset comparison.
@@
    Standardized Benchmarking Environment. We introduce a standardized evaluation framework, PoseBench3D, that consolidates four commonly used human pose estimation datasets—H36M[10], GPA[34], 3DPW[21], and SURREAL[32]—while providing modularity for future datasets. Our framework simplifies cross-dataset evaluation by requiring, simply, a single user-supplied configuration file (cf. section A.1) specifying model, experiment, and user customization details.
@@ 2 Related Work
Current Cross-Dataset Evaluation. While several methods have attempted to address generalization, none have demonstrated success beyond testing on a single additional dataset beyond their training set. In particular, TCPFormer [14], FinePOSE [37], MotionBERT [48], and MotionAGFormer [28] all train on Human3.6M [10] and test only on 3DPW [21]. Although each introduces novelty—TCPFormer in lifting, FinePOSE in architectural design, and MotionBERT and MotionAGFormer in binary decisions regarding the nature of the test image—none have demonstrated the ability to generalize across multiple datasets, let alone real-world scenarios. In contrast, PoseBench3D automates cross-dataset evaluation, eliminating long-existing friction in benchmarking across datasets and addressing a key limitation that has held back progress in the field.
Shortage of Testing Environments. Although we commend AdaptPose [7] for evaluating its fine-tuned version of [24] and [8] across four datasets—Human3.6M, 3DHP, 3DPW, and Ski-Pose—they do not provide a standardized evaluation framework for others to test their own models. Wang et al. [35], Manzur et al. [20], and Gong et al. [8] also report cross-dataset results, which aligns more closely with our goals. Nevertheless, there remains a lack of standardized cross-dataset evaluation frameworks that enable consistent benchmarking across multiple works.
@@ 3 Overview of PoseBench3D
Our framework, illustrated in the high-level overview in Figure 4, is designed to support the evaluation of 2D-to-3D human pose lifting networks. In what follows, we provide a brief summary of the supported 2D-to-3D pose estimation datasets, outline the curated set of benchmark models included in our evaluation, and describe the techniques employed to standardize these datasets under a unified and fair comparison framework.
@@ 3.1 Benchmark Datasets
Our framework currently supports four datasets for evaluation, depicted in table 1. The Human3.6M dataset (H36M) [10] consists of over 3.6 million annotated 3D human poses from 15 natural activity scenarios. This dataset is currently the most popular for use in the field of 3D human pose estimation, but its controlled lab environment limits its generalization potential. Geometric Pose Affordance (GPA) [34] focuses on the interaction between humans and objects. With 305 thousand RGB images, GPA emphasizes environmental interaction. The 3D Poses in the Wild (3DPW) dataset [21] emphasizes real-world, unconstrained environments with pose data captured using IMU sensors along with automatically generated 2D-to-3D pose associations. SURREAL [32] is a large synthetic dataset for 3D pose estimation, derived from the CMU Motion Capture database [5]. It uses MoSh [19] to match SMPL [18] parameters to raw marker paths. Since the dataset is synthetic and uses human skeletal models, both 2D and 3D ground truth joint locations are available to arbitrary precision.
@@ 3.2 Supported Model Architectures
Our framework is set-up to support Pose Lifting Networks models out-of-the-box, with the goal of supporting video- and image-based models in future work. 2D-to-3D pose estimation models, a.k.a. “Lifting Networks”, are widely used for their flexibility, augmentation capabilities, and ease of implementation. Given that 2D keypoint-based representations lack many visual cues present in images (such as occlusion), it simplifies the task of 3D pose estimation–often at the cost of generalization. As reported by [35, 20], they show poor performance when tested against other datasets. Therefore, we survey the lifting networks reported by previous works [22, 42, 44, 39, 41, 1, 6, 8, 45, 43, 40, 4, 49] and pick 18 configurations based on their availability and adaptability. Gong et al. [8] reported results on Martinez et al. [22], Zhao et al. [42], Pavllo et al. [25], and Cai et al. [1] with optimized and un-optimized variants tested on the H36M and 3DPW datasets. Since the checkpoints are already provided, we used these checkpoints to re-evaluate the models on all four datasets with our framework. We also gathered five transformer-based models [45, 43, 44, 26, 13], two graph convolution models [39, 42], a spatio-temporal encoder model [41], and a diffusion-based model [27]. Given that various models from the authors work were missing checkpoints or were trained on 17 joint skeleton configurations, we decided to standardize testing by retraining these models from scratch on the more common 16 joint skeleton configuration to remain fair and consistent across comparisons. We further ensure the retrained model is as good as the original one by comparing the reported same-dataset results directly.
@@ 3.3 Dataset Preparation
To unify four distinct datasets under a common processing interface, we first identified the core elements required by lifting-based 3D pose estimation networks. These models typically take 2D keypoints as input and predict 3D joint coordinates as output. The 3D pose is usually expressed either in world-space or camera-space, with the latter being both statistically justified and widely adopted due to its alignment with image-space supervision. Additionally, most models standardize the pose by centering the skeleton at the hip joint. As shown in table 1 and discussed in [35], datasets differ significantly in terms of joint definitions, camera intrinsics/extrinsics, image resolution, and other parameters. As such, each dataset requires tailored preprocessing pipelines. For efficiency, we preprocessed and cached all dataset files as NumPy zipped archives, enabling faster data loading. Although different datasets call for different preprocessing techniques, the most important operation is acquiring and converting 3D joint positions into camera coordinates. Once this is established, projecting to 2D space follows a standardized procedure.
@@
The H36M [10] dataset provides four cameras with their intrinsic and extrinsic matrices. While the intrinsic matrices stay the same for all 15 action sequences across 11 subjects, the extrinsic parameters (e.g. orientation and translation) change. There are 15 actions, 2 subactions per action, and 4 cameras, all acted by 11 subjects. Since the 3D pose estimation task focuses training on subjects 1, 5, 6, 7, and 8 and testing on subjects 9 and 11, we only provide data points for these subjects. To avoid redundancy, we only sample frames that have noticeable change in the pose. The world-space 3D coordinates ($\mathrm{X}\in\mathbb{R}^{N\times 3}$) are then converted into camera-space coordinates with the help of $(\mathrm{R}\cdot(\mathrm{X}^{\top}-\mathrm{t}))^{\top}$, where $\mathrm{R}\in\mathbb{R}^{3\times 3}$ is the rotation matrix and $\mathrm{t}$ is the translation vector. The 2D points were projected using the perspective projection equation to obtain the ground-truth 2D keypoints. On the other hand, the GPA [34] dataset provides intrinsic and extrinsic parameters for every frame where all the annotations are provided in a json file. The key difference in processing the dataset was the camera-space coordinate transformation. We use $\mathrm{R_{i}}^{\top}\cdot\mathrm{X}_{i}^{\top}+\mathrm{t_{i}}$, where the notification follows a similar convention to that of the H36M dataset. The translation vector provided was in centimeters in the original dataset, where the 3D coordinates were found to be in millimeters. The rotation matrix was also stored in a vector format. The 3DPW dataset [21] immediately differentiates itself from the rest of the datasets in terms of formatting. The dataset provides camera extrinsics and intrinsics per sequence, and each sequence can contain multiple subjects. We process each subject separately. Since the camera parameters are provided as an extrinsic matrix ($\mathbb{R}^{4\times 4}$), a multiplication with the world space coordinates expressed as homogenous coordinates is enough to obtain the camera space coordinates. The SURREAL dataset [32] required minimal preprocessing for camera coordinate conversion. However, we identified several invalid data batches, which consistently caused large pose estimation errors. To maintain consistency in evaluation, we excluded these anomalous samples, which accounted for less than 1% of the dataset.
@@ 3.4 Framework Initialization
[…]
@@ 4.2 Cross-Dataset Evaluation
We report the MPJPE scores in table 2 and PA-MPJPE scores in table 3. We also include two cross-dataset results available from prior work marked with $\ddagger$ in the tables.
@@
Table 2: Cross-dataset evaluation sorted by decreasing average MPJPE (mm). All models are trained on H36M. Lower is better. $\dagger$: PoseAug-optimized; $\diamond$: unoptimized variant; $\bullet$: retrained from scratch; $\ddagger$: reported from prior work.
@@
Model Name H36M GPA 3DPW SURREAL Average Error (MPJPE $\downarrow$)
@@
Table 3: Cross-dataset evaluation sorted by decreasing average PA-MPJPE (mm). All models are trained on H36M. Lower is better. $\dagger$: PoseAug-optimized; $\diamond$: unoptimized variant; $\bullet$: retrained from scratch; $\ddagger$: reported from prior work.
@@
Model Name H36M GPA 3DPW SURREAL Average Error (PA-MPJPE $\downarrow$)
@@
We also analyze per-joint position errors under both protocols in fig. 5, averaged across all 18 model configurations and all four datasets. While the endpoints of the human body—such as the wrists, ankles, and head—are more susceptible to errors, note that Procrustes alignment alone results in a substantial reduction in these errors (highlighted in red). Since Procrustes alignment accounts for rigid transformations, this highlights that human “pose” estimation is distinct from “position” estimation. Therefore, future efforts should focus on reducing angular errors in the bones rather than solely minimizing positional discrepancies.

### v2

@@ Abstract
Reliable three-dimensional human pose estimation (3D HPE) remains challenging due to the differences in viewpoints, environments, and camera conventions among datasets. As a result, methods that achieve near-optimal in-dataset accuracy often degrade on unseen datasets. In practice, however, systems must adapt to diverse viewpoints, environments, and camera setups—conditions that differ significantly from those encountered during training, which is often the case in real-world scenarios. Measuring cross-dataset performance is a vital process, but extremely labor-intensive when done manually for human pose estimation. To address these challenges, we automate this evaluation using PoseBench3D, a standardized testing framework that enables consistent and fair cross-dataset comparisons on previously unseen data. PoseBench3D streamlines testing across four widely used 3D HPE datasets via a single, configurable interface. Using this framework, we re-evaluate 18 methods and report over 100 cross-dataset results under Protocol 1: MPJPE and Protocol 2: PA-MPJPE, revealing systematic generalization gaps and the impact of common preprocessing and dataset setup choices. The PoseBench3D code is found at: https://github.com/bryanjvela/PoseBench3D.
@@ 1 Introduction
Three-dimensional Human Pose Estimation (HPE) has gathered substantial interest for its critical role in applications such as healthcare [37, 29], action recognition [17, 18], military operations [47, 13], human-computer interaction [30, 39], and virtual/augmented reality [23]—among many others. Despite remarkable progress in recent years [15, 31, 48, 28, 10, 34], most work focuses on performance within a single, controlled dataset. The majority of existing work adopts either a direct-from-image approach or a two-stage approach, with the latter being widely used for its flexibility. In the two-stage approach, the first stage detects 2D keypoints, and the second stage, known as the “lifting network” [22, 43, 16, 5, 49, 12, 41] maps the 2D keypoints to 3D pose estimations. However, these two-stage models show poor generalization across various different datasets [36, 21].
@@
To assess the generalization of these models, we consider cross-dataset performance: the task of training on one dataset and testing on another. Machine learning models are generally susceptible to overfitting to the bias and variance of the data distribution they were trained on; this is particularly pronounced in 3D human pose estimation datasets due to variations in data collection procedures, where the underlying distributions of datasets often differ significantly. For example, the H36M dataset [11] was captured using a four-camera setup, while GPA [35] used five, and both 3DPW [33] and SURREAL [32] relied on a single-camera configuration. In Fig.˜1, we show the frequency histogram of camera positions relative to the subject in spherical coordinates. While elevation distributions are largely consistent across datasets (cf. Fig.˜1(a)), the azimuthal angles vary significantly (cf. Fig.˜1(b)).
Beyond camera setup, joint placement conventions also differ between datasets. For instance, there is currently no standardized convention for placing the hip or spine joints (cf. Fig.˜2)—note the difference in width of hips in Fig.˜2(a) and Fig.˜2(c). Additionally, since datasets are collected with varying action sets, the captured range of motion also varies. In Fig.˜3, the blue dots represent the recorded range of motion, whereas the yellow region shows the valid range. Since these two datasets do not explore the entire valid range of motion, normalization performed with one will not translate well onto the other.
Research Gap. These differences mean that, although any method may achieve near-optimal performance within its own dataset, it is likely to perform poorly across differing datasets. Measuring cross-dataset performance is paramount, but is extremely labor-intensive to perform manually. Our goal in this paper is to automate the task of cross-dataset comparison.
@@
    Standardized Benchmarking Environment. We introduce a standardized evaluation framework, PoseBench3D, that consolidates four commonly used human pose estimation datasets—H36M[11], GPA[35], 3DPW[33], and SURREAL[32]—while providing modularity for future datasets. Our framework simplifies cross-dataset evaluation by requiring, simply, a single user-supplied configuration file specifying model, experiment, and user customization details.
@@ 2 Related Work
Current Cross‑Dataset Evaluation. While several methods have attempted to address generalization, none have demonstrated success beyond testing on a single additional dataset beyond their training set. In particular, TCPFormer [15], FinePOSE [38], and MotionAGFormer [28] all train on Human3.6M [11] and test only on 3DPW [33]. Although each introduces novelty—TCPFormer in lifting, FinePOSE in architectural design, and MotionBERT and MotionAGFormer in binary decisions regarding the nature of the test image—none have demonstrated the ability to generalize across multiple datasets, let alone real-world scenarios. In contrast, PoseBench3D automates cross-dataset evaluation, eliminating long-existing friction in benchmarking across datasets and addressing a key limitation that has held back progress in the field.
Shortage of Testing Environments. Although we commend AdaptPose [8] for evaluating its fine-tuned version of [24] and [9] across four datasets—Human3.6M, 3DHP, 3DPW, and Ski‑Pose—they do not provide a standardized evaluation framework for others to test their own models. Wang et al. [36], Manzur et al. [21], and Gong et al. [9] also report cross-dataset results, which aligns more closely with our goals. Nevertheless, there remains a lack of standardized cross-dataset evaluation frameworks that enable consistent benchmarking across multiple works.
@@ 3 Overview of PoseBench3D
Our framework, illustrated in the high-level overview in Figure˜4, is designed to support the evaluation of 2D-to-3D human pose lifting networks. In what follows, we provide a brief summary of the supported 2D-to-3D pose estimation datasets, outline the curated set of benchmark models included in our evaluation, and describe the techniques employed to standardize these datasets under a unified and fair comparison framework.
@@ 3.1 Benchmark Datasets
Our framework currently supports four datasets for evaluation, depicted in Tab.˜1. The Human3.6M dataset (H36M) [11] consists of over 3.6 million annotated 3D human poses from 15 natural activity scenarios. This dataset is currently the most popular for use in the field of 3D human pose estimation, but its controlled lab environment limits its generalization potential. Geometric Pose Affordance (GPA) [35] focuses on the interaction between humans and objects. With $\approx$ 305,000 RGB images, GPA emphasizes environmental interaction. The 3D Poses in the Wild (3DPW) dataset [33] emphasizes real-world, unconstrained environments with pose data captured using IMU sensors along with automatically generated 2D-to-3D pose associations. SURREAL [32] is a large synthetic dataset for 3D pose estimation, derived from the CMU Motion Capture database [6]. It uses MoSh [20] to match SMPL [19] parameters to raw marker paths. Since the dataset is synthetic and uses human skeletal models, both 2D and 3D ground truth joint locations are available to arbitrary precision.
@@ 3.2 Supported Model Architectures
Our framework is set-up to support Pose Lifting Networks models out-of-the-box, with the goal of supporting video- and image-based models in future work. 2D-to-3D pose estimation models, a.k.a. “Lifting Networks”, are widely used for their flexibility, augmentation capabilities, and ease of implementation. Given that 2D keypoint-based representations lack many visual cues present in images (such as occlusion), it simplifies the task of 3D pose estimation–often at the cost of generalization [36, 21]. Therefore, we survey the lifting networks reported by previous works [22, 43, 45, 40, 42, 2, 7, 9, 46, 44, 41, 5, 49] and pick 18 configurations based on their availability and adaptability. Gong et al. [9] reported results on Martinez et al. [22], Zhao et al. [43], Pavllo et al. [25], and Cai et al. [2] with optimized and un-optimized variants tested on the H36M and 3DPW datasets. Since the checkpoints are already provided, we used these checkpoints to re-evaluate the models on all four datasets with our framework. We also gathered five transformer-based models [46, 44, 45, 26, 14], two graph convolution models [40, 43], a spatio-temporal encoder model [42], and a diffusion-based model [27]. Given that some of these works omitted some checkpoints or were trained on 17 joint skeleton configurations, we decided to standardize testing by retraining these models from scratch on the more common 16 joint skeleton configuration to remain fair and consistent across comparisons. We further ensure the retrained model is as good as the original one by comparing the reported same-dataset results directly.
@@ 3.3 Dataset Preparation
To unify four distinct datasets under a common processing interface, we first identified the core elements required by lifting-based 3D pose estimation networks. These models typically take 2D keypoints as input and predict 3D joint coordinates as output. The 3D pose is usually expressed either in world-space or camera-space, with the latter being both statistically justified and widely adopted due to its alignment with image-space supervision. Additionally, most models standardize the pose by centering the skeleton at the hip joint. As shown in Tab.˜1 and discussed in [36], datasets differ significantly in terms of joint definitions, camera intrinsics/extrinsics, image resolution, and other parameters. As such, each dataset requires tailored preprocessing pipelines. For efficiency, we preprocessed and cached all dataset files as NumPy zipped archives, enabling faster data loading. Although different datasets call for different preprocessing techniques, the most important operation is acquiring and converting 3D joint positions into camera coordinates. Once this is established, projecting to 2D space follows a standardized procedure.
@@
The H36M [11] dataset provides four cameras with their intrinsic and extrinsic matrices. While the intrinsic matrices stay the same for all 15 action sequences across 11 subjects, the extrinsic parameters (e.g., orientation and translation) change. There are 15 actions, 2 subactions per action, and 4 cameras, all acted by 11 subjects. Since the 3D pose estimation task focuses training on subjects 1, 5, 6, 7, and 8 and testing on subjects 9 and 11, we only provide data points for these subjects. To avoid redundancy, we only sample frames that have noticeable change in the pose. The world-space 3D coordinates ($\mathrm{X}\in\mathbb{R}^{N\times 3}$) are then converted into camera-space coordinates with the help of $(\mathrm{R}\cdot(\mathrm{X}^{\top}-\mathrm{t}))^{\top}$, where $\mathrm{R}\in\mathbb{R}^{3\times 3}$ is the rotation matrix and $\mathrm{t}$ is the translation vector. The 2D points were projected using the perspective projection equation to obtain the ground-truth 2D keypoints. On the other hand, the GPA [35] dataset provides intrinsic and extrinsic parameters for every frame where all the annotations are provided in a json file. The key difference in processing the dataset was the camera-space coordinate transformation. We use $\mathrm{R_{i}}^{\top}\cdot\mathrm{X}_{i}^{\top}+\mathrm{t_{i}}$, where the notification follows a similar convention to that of the H36M dataset. The translation vector provided was in centimeters in the original dataset, where the 3D coordinates were found to be in millimeters. The rotation matrix was also stored in a vector format. The 3DPW dataset [33] immediately differentiates itself from the rest of the datasets in terms of formatting. The dataset provides camera extrinsics and intrinsics per sequence, and each sequence can contain multiple subjects. We process each subject separately. Since the camera parameters are provided as an extrinsic matrix ($\mathbb{R}^{4\times 4}$), a multiplication with the world space coordinates expressed as homogenous coordinates is enough to obtain the camera space coordinates. The SURREAL dataset [32] required minimal preprocessing for camera coordinate conversion. However, we identified several invalid data batches, which consistently caused large pose estimation errors. To maintain consistency in evaluation, we excluded these anomalous samples, which accounted for less than 1% of the dataset.
@@ 3.4 Framework Initialization
Our framework consists of five key modules, as illustrated in Fig.˜4. Experiments are orchestrated through a global configuration file in YAML format, specifying critical parameters such as model_type, num_workers, trained_on_normalized_data, output_3d, video_mode, num_joints, and num_frames, among others. The framework supports model checkpoints saved in either JIT or ONNX format. The trained_on_normalized_data flag indicates whether the model was trained on normalized data, while the video_mode toggle enables inference on multiple frames simultaneously, which is essential for temporal models. By adopting this configuration-driven approach, models can be seamlessly loaded as abstract entities, simplifying the complexity associated with conducting diverse experiments. Specifically, the Model Module initializes the chosen model, and the Dataset Module processes and prepares the datasets according to the configuration settings. The selection of joints used by the Dataset Module is determined by the num_joints parameter provided in the configuration. To maintain consistency across experiments, we primarily employ a standardized 16-joint ordering: hip, right hip, right knee, right ankle, left hip, left knee, left ankle, spine, neck, head, left shoulder, left elbow, left wrist, right shoulder, right elbow, and right wrist. For certain experiments, we alternatively employ a 14-joint subset, excluding the head and spine. All evaluated models require either screen-space normalization or z-score standardization. Screen-space normalization scales pixel coordinates from the original image dimensions $(0,w)$ (width) and $(0,h)$ (height) to a unit interval $(0,1)$. In contrast, z-score standardization involves normalizing joint coordinates, $X$ with $\frac{X-\mu}{\sigma}$, where $\mu$ is the mean and $\sigma$ is the standard deviation.
@@ 4.2 Cross-Dataset Evaluation
We report the MPJPE scores in Tab.˜2 and PA-MPJPE scores in Tab.˜3. We also include two cross-dataset results available from prior work marked with $\ddagger$ in the tables.
@@
Table 2: Cross-dataset evaluation sorted by decreasing average MPJPE ($m$). All models are trained on H36M. Lower is better. $\dagger$: PoseAug-optimized; $\diamond$: unoptimized variant; $\bullet$: retrained from scratch; $\ddagger$: reported from prior work.
@@
Model Name Human3.6M GPA 3DPW SURREAL Average Error (MPJPE $\downarrow$)
@@
Table 3: Cross-dataset evaluation sorted by decreasing average PA-MPJPE ($m$). All models are trained on H36M. Lower is better. $\dagger$: PoseAug-optimized; $\diamond$: unoptimized variant; $\bullet$: retrained from scratch; $\ddagger$: reported from prior work.
@@
Model Name Human3.6M GPA 3DPW SURREAL Average Error (PA-MPJPE $\downarrow$)
@@
We also analyze per-joint position errors under both protocols in Fig.˜5, averaged across all 18 model configurations and all four datasets. While the endpoints of the human body—such as the wrists, ankles, and head—are more susceptible to errors, note that Procrustes alignment alone results in a substantial reduction in these errors (highlighted in red). Since Procrustes alignment accounts for rigid transformations, this highlights that human “pose” estimation is distinct from “position” estimation. Therefore, future efforts should focus on reducing angular errors in the bones rather than solely minimizing positional discrepancies.